# Symplectic Quantization: numerical results for the Feynman propagator on a 1+1 lattice and the theoretical relation with Quantum Field Theory

### v2

@@ I Introduction
Since its invention by Kenneth Wilson [1], lattice field theory had an enormous development [2, 3] as a method to handle non-perturbative problems in quantum field theory, in particular concerning the theory of strong interactions with respecto to problems such as the estimate of hadronic masses [4] or heavy ions collisions [5]. Nevertheless, despite the great achievements, any numerical approach to quantum field theory on the lattice has retained so far a major limitation: all importance-sampling protocols are well defined only for Euclidean field theory. Since the latter is obtained by Wick-rotating real into imaginary time, the causal structure of correlation functions in quantum field theory cannot be usually sampled numerically. The convenience/necessity to analytically continue real to immaginary time is to transform the Feynman path integral, characterized by the oscillating factor $\exp(iS[\phi]/\hbar)$  —  $S[\phi]$ is the relativistic action and $\phi$ a generic quantum field  — into a normalizable probability density $\exp(-S_{E}[\phi]/\hbar)$, with $\hbar$ playing the same role of temperature in the Boltzmann weight of statistical mechanics and where $S_{E}[\phi]$ is the positive-definite Euclidean action. The mapping to imaginary time has been so far the unavoidable condition to set up any importance sampling numerical protocol to study the quantum fluctuations of fields. The goal of the present work is to present a conceptual framework and a method which goes beyond this limitation, allowing for a straightforward procedure to numerically sample quantum fluctuations of fields in Minkowskian spacetime. The use of imaginary time and Euclidean field theory forbids the representation on the lattice of any process or phenomenon intrinsically related to the causal structure of space-time, in particular all processes on the light cone. By definition the probability density $\exp(-S_{E}[\phi]/\hbar)$ works as an effective “equilibrium” measure for quantum fluctuations. Any importance-sampling protocol built from the Euclidean weight projects, for large lattice sizes, on the ground states of the corresponding Minkowskian theory. In fact, Monte Carlo simulations built from Euclidean field theory allow us to reproduce with extreme precision the physics of stable/equilibrium bound states of the strong interactions [4], whereas it has been so far much more problematic to reproduce metastable resonances with short lifetimes, like for instance tetraquark or pentaquark states [6, 7], or the dynamics of scattering processes with a strong relativistic character, namely processes with a different number of degrees of freedom in the asymptotic initial and final states. It is for this reason that we believe it is of crucial interest the possibility to test numerically any new proposal for a quantum field theory formulation which allows first to define and then to study the dynamics of quantum fields fluctuations directly in Minkowski space-time.
@@
An interesting idea in this direction, namely the proposal of a functional approach to field theory which is well defined from the probabilistic point of view already in Lorentzian space-time, has been recently put forward by one of us and goes under the name of “symplectic quantization” [8, 9, 10]. The first ingredient of this approach is to assume, for a given quantum field $\phi(x)$, with $x=(ct,{\bf x})$ a point in four-dimensional space-time, a dependence on an additional time parameter $\tau$:
$\phi(x)\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ \phi(x,\tau),$ (1)
@@
which controls the continuous sequence of quantum fluctuations in each point of space-time. Theories with such an additional time parameter are not a novelty, the whole Parisi-Wu stochastic quantization approach being based on this idea [11, 12]: it is for the reason that we deem as a necessary prelude, before presenting symplectic quantization, to summarize the main features of stochastic quantization. The very simple idea of stochastic quantization is to introduce, as we said, a fictitious dynamics for the quantum fluctuations of an Euclidean field theory. This dynamics, following the analogy with statistical mechanics, is the one of a Langevin equation for the field $\phi(x,\tau)$:
@@
where $S_{E}[\phi]$ is the Euclidean action and $\eta(x,\tau)$ is a “white noise”, that is a stochastic process characterized by its first moments
$\langle\eta(x,\tau)\rangle=0$
$\langle\eta(x,\tau)\eta(y,\tau^{\prime})\rangle=2\leavevmode\nobreak\ \hbar\leavevmode\nobreak\ \delta(\tau-\tau^{\prime})\delta^{4}(x-y).$ (3)
The Euclidean weight of field fluctuations is obtained via the Langevin equation in the following manner. It is a proved fact that a non-linear stochastic equation for the field such as Eq. (2) is related to a linear equation for the probability distribution of the field, the Fokker-Planck equation:
$\frac{d}{d\tau}P_{\tau}(\phi)$ $=\left[\frac{\delta^{2}}{\delta\phi(x)^{2}}+\frac{\delta}{\delta\phi(x)}\cdot S_{E}[\phi]\right]P_{\tau}(\phi)$
$=L_{\text{FP}}[\phi]\circ P_{\tau}(\phi).$ (4)
The link between Fokker-Planck and Euclidean field theory comes with the fact that Eq. (4) admits as a stationary solution the equilibrium distribution $P_{\text{eq}}(\phi)$:
$P_{\text{eq}}(\phi)=\frac{1}{\mathcal{Z}}\exp\left(-\frac{1}{\hbar}S_{E}[\phi]\right)$ (5)
Within the stochastic quantization approach, Euclidean multipoint correlation functions are therefore obtained as the infinite-time limit of equal-time multipoint correlations computed along the stochastic dynamics, in such a way that any particular realization of the noise $\eta(x,\tau)$ has no influence on the final result. The presence of the additional time $\tau$ therefore did not stimulate any conceptual discussion on its interpretation, being it regarded just a sort of “computational trick”. Therefore, while on the one hand stochastic quantization had the merit of introducing the key idea of an additional time controlling the sequence of quantum fluctuations, often referred to in the literature as “fictitious time” [11], on the other hand this proposal did not trigger a substantial conceptual advance on the foundations of quantum field theory, being such time just an auxiliary variable and being all transients in the dynamics regarded just as unphysical.
Now, before introducing explicitly the symplectic quantization formalism let us mention, among the several spin-off of stochastic quantization which were proposed in the past decades, one which somehow preluded our proposal: the microcanonical approach to quantum field theory introduced in [13]. The crucial novelty of this generalized Hamiltonian dynamics, introduced to overcome the numerical difficulties in the implementation of lattice gauge theory with fermions [13], was the presence of conjugated momenta proportional to the rate of variation of fields with respect to $\tau$. On the basis of the equivalence between the Euclidean measure of Eq. (5) and a corresponding microcanonical one, it was noted that a deterministic dynamics of the kind
$\frac{d^{2}\phi}{d\tau^{2}}=-\frac{\delta S_{E}[\phi]}{\delta\phi(x,\tau)}$ (6)
was leading to the same asymptotic probability distribution of fields produced by the stochastic dynamics in Eq. (2), with the major computational advantage of being more suited to parallel updates of the variables, something particularly useful in the case of the non-local bosonic actions obtained from the integration of fermionic variables [3]. The introduction of canonical conjugated momenta of the quantum fields with respect to the intrinsic time was thoroughly discussed in [14], also in this case within the Euclidean field theory framework.
Let us now present the main novelty of symplectic quantization, as announced in [8, 9]: this is the idea that the correct numerical sampling of quantum fluctuation in a Lorentzian space-time can be done directly only in a sort of generalized microcanonical ensemble, built by adding the conjugated momenta with respect to the intrinsic time $\tau$ directly to the relativistic action with Minkowski signature, with no need of considering any sort of analytic continuation from real to immaginary time and therefore preserving the causal structure of space-time. This intuition has been strongly inspired from the evidence that in statistical mechanics there are physically important situations where only the microcanonical ensemble is well defined [15, 16], namely physical phenomena which cannot be described within the canonical ensemble.
In the present work we show how to sample the quantum fluctuations of a real scalar $\lambda\phi^{4}$ theory in 1+1 Lorentzian space-time dimensions by studying a generalized Hamiltonian dynamics in the intrinsic time $\tau$: our main result is the unprecedented numerical measure of a two-point correlation function which, for small values of the interaction constant $\lambda$, reproduces qualitatively well the space-time structure of the free Feynman propagator. Despite this success in capturing the causal structure of space-time in numerical simulations, a task which has been so far out of reach for any numerical approach based on Euclidean field theory, symplectic quantizaton, at least in the form presented here, does not precisely correspond to a sampling of fluctuations in the original quantum field theory (although accomplishing tasks precluded to Euclidean field theory). The precise relation between symplectic quantization in its present form and quantum field theory is discussed in the second part of this paper, where we present a direct calculation of the microcanonical partition function in the large-$M$ limit, where $M$ is the number of degrees of freedom in the lattice. The result of this calculation is to show that the sampling of quantum fluctuation with our symplectic dynamics is equivalent to the sampling according to a canonical weight of the kind
@@
where the results of our simulations correspond to fixing $z=1$ in Eq. (7). Clearly, a probability density like the one in Eq. (7) is not equivalent to the complex amplitude which enters the Feynman path integral, but can be related to it by analytic continuation in the complex $z$ plane. Investigations in this direction are in progress and will be presented in a forthcoming paper explaining how the symplectic quantization approach works for a quantum particle in a harmonic potential, with particular emphasis on how the analytic continuation to complex $z$ plane of the above density, Eq. (7), must be handled [17].
II Symplectic Quantization: from dynamics to ensemble averages
Let us summarize here the main steps for the derivation of the symplectic quantization dynamics. First of all, inspired by the stochastic quantization approach [11, 12], we assume that quantum fields $\phi(x,\tau)$ depend on an additional time variable $\tau$ which parametrizes the dynamics of quantum fluctuations in a given point of Minkowski space-time. Since for a relativistic quantum field theory the ambient space includes observer’s time, necessarily the intrinsic time $\tau$ must be a different variable, as thoroughly discussed in [8, 9]. The symplectic quantization approach to field theory assumes, consistently with the existence of an intrinsic time $\tau$, the existence of conjugated momenta of the kind
@@ II Symplectic Quantization: from dynamics to ensemble averages
$\mathbb{L}[\phi,\dot{\phi}]=\int d^{d}x\leavevmode\nobreak\ \left[\frac{1}{2c_{s}^{2}}\dot{\phi}^{2}(x)+S[\phi]\right],$ (9)
@@
$\mathbb{H}[\phi,\pi]=\frac{1}{2}\int d^{d}x\leavevmode\nobreak\ c_{s}^{2}\leavevmode\nobreak\ \pi^{2}(x)-S[\phi]$
@@
$=\int d^{d}x\leavevmode\nobreak\ \left[\frac{c_{s}^{2}}{2}\pi^{2}(x)+\frac{1}{2}\phi\leavevmode\nobreak\ \partial_{0}^{2}\phi-\sum_{i=1}^{d}\phi\leavevmode\nobreak\ \partial_{i}^{2}\phi+V[\phi]\right]$ (12)
@@
$\mathbb{H}[\phi,\pi]$ $=\mathbb{K}[\pi]+\mathbb{V}[\phi]$
@@
$\mathbb{K}[\pi]$ $=\frac{1}{2}\int d^{d}x\leavevmode\nobreak\ \pi^{2}(x).$ (16)
@@
Having defined the above deterministic dynamics for the quantum fluctuations of the field $\phi(x,\tau)$, Eq. (14), one can then legitimately wonder how this functional formalism connects to the standard one, for instance to the standard Feynman path-integral formulation of quantum field theory. In first instance, in order to find out the connection between the dynamic approach of symplectic quantization and any functional formulation of field theory where the additional time $\tau$ is absent, one needs to find out which probability density $\rho[\phi(x)]$ corresponds to the dynamics. There are then two possibilities: $\rho[\phi(x)]$ might have either an integral or a punctual correspondence with quantum field theory. The integral correspondence is when the path-integral is recovered by means of an integral transformation, as it will be discussed below here, and the punctual correspondence is when, in a certain limit, one can directly show that $\rho[\phi(x)]\approx e^{iS[\phi]/\hbar}$. This last procedure is more subtle, because it requires an analytic continuation of the action and of its degrees of freedom in the complex plane, and will be discussed thoroughly in [17]. But let us now go back to the relation between symplectic dynamics and probability densities, for which we make an ergodic hypothesis for the Hamiltonian dynamics in Eq. (13): if we assume that this dynamics samples at long time $\tau$ the constant generalized energy hypersurface with uniform probability [8, 9], then we can associate to the dynamics of Eq. (13) the following measure:
$\rho_{\text{micro}}[\phi(x)]=\frac{1}{\Omega[\mathcal{A}]}\leavevmode\nobreak\ \delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]\right),$ (18)
@@
$\Omega[\mathcal{A}]=\int\mathcal{D}\phi\mathcal{D}\pi\leavevmode\nobreak\ \delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]\right),$ (19)
@@
[…]
where $\tau_{0}$ is a large enough time for the system to have reached stationarity and “lost memory” of initial conditions. How to relate then the microcanonical partition function in Eq. (19) to the path integral? It is quite intuitive to understand that the two expression must be related by some sort of statistical ensemble change. The crucial point of this change of ensemble, as stressed already in [8], is that the microcanonical partition function $\Omega[\mathcal{A}]$ is built on the conservation of a non-positive quantity, the generalized Hamiltonian $\mathbb{H}[\phi,\pi]$. The latter, from the point of view of physical dimensions, is a relativistic action: it therefore takes both arbitrarily large positive and arbitrarily large negative values due to the negative sign in front of the coordinate-time derivative term in the second line of Eq. (12). The absence of positive definiteness for the generalized Hamiltonian $\mathbb{H}[\phi,\pi]$, which is the true relativistic signature of the theory, is what forbids a standard change of ensemble with a Laplace transform, that is customary in statistical mechanics when passing from microcanonical to canonical. The only integral transform which allows us to map formally the ensemble where $\mathbb{H}[\phi,\pi]$ is constrained to the one where it is free to fluctuate is the Fourier transform. It is by Fourier transforming the microcanonical partition function $\Omega[\mathcal{A}]$ that one obtains straightforwardly the Feynman path integral:
[…]
$=\int\mathcal{D}\phi\leavevmode\nobreak\ \mathcal{D}\pi\leavevmode\nobreak\ e^{-\frac{i}{2}u\int d^{d}x\leavevmode\nobreak\ \pi^{2}(x)+iuS[\phi]}$
$=\mathcal{N}(u)\int\mathcal{D}\phi\leavevmode\nobreak\ e^{iuS[\phi]},$ (22)
@@
[…]
The one above is to our knowledge the first derivation from first principles of the Feynman path-integral formula in the context of a more extended framework. We could say that this larger framework is the statistical mechanics of action-preserving systems, opposed to the statistical mechanics of energy-preserving systems, which is the standard one. And it is precisely the fact that statistical ensembles are built on the conservation of a non positive-defined quantity, which is the true landmark of the relativistic nature of the theory, that determines the fact that locally we can only access complex probability amplitudes and not real probabilities. From the perspective of symplectic quantization the replacement at the local level of probabilities with probability amplitudes is therefore a direct consequence of special relativity and a wise use of statistical ensembles. To better understand this statement let us consider an unrealistic situation where the symplectic action (generalized Hamiltonian) $\mathbb{H}[\phi,\pi]$ was positive definite. In this case one could change ensemble with Laplace rather than Fourier transform,
[…]
@@
We have just shown how the standard path-integral formulation can be recovered, on the basis of an ergodicity assumption, from the symplectic quantization dynamics approach and which is the role played by $\hbar$ within this, let us say, change of ensemble. At the same time it is not only legitimate but also necessary to wonder if and how there is a quantization constraint involving $\hbar$ which can be imposed directly on the microcanonical ensemble of symplectic quantization. The indication coming from the stochastic quantization framework is that $\hbar$ must play a role analogous to that of temperature. Therefore, as suggested in [10], we believe that the most natural assumption for the role of $\hbar$ in the symplectic quantization formalism is to be analogous to the microcanonical temperature:
@@
$\langle\pi(x)\pi(y)\rangle=\frac{\hbar}{2}\leavevmode\nobreak\ \delta^{(d)}(x-y),$ (26)
@@
[…]
@@
$\langle\pi(x_{i})\pi(x_{j})\rangle=\frac{\hbar}{2}\leavevmode\nobreak\ \frac{\delta_{ij}}{a^{d}},$ (28)

### v3

@@ 1 Introduction
Since its invention by Kenneth Wilson W74 , lattice field theory had an enormous development C83 ; MM94 as a method to handle non-perturbative problems in quantum field theory, in particular concerning the theory of strong interactions with respecto to problems such as the estimate of hadronic masses FH12 or heavy ions collisions R18 . Nevertheless, despite the great achievements, any numerical approach to quantum field theory on the lattice has retained so far a major limitation: all importance-sampling protocols are well defined only for Euclidean field theory. Since the latter is obtained by Wick-rotating real into imaginary time, the causal structure of correlation functions in quantum field theory cannot be usually sampled numerically. The convenience/necessity to analytically continue real to immaginary time is to transform the Feynman path integral, characterized by the oscillating factor $\exp(iS[\phi]/\hbar)$  —  $S[\phi]$ is the relativistic action and $\phi$ a generic quantum field  — into a normalizable probability density $\exp(-S_{E}[\phi]/\hbar)$, with $\hbar$ playing the same role of temperature in the Boltzmann weight of statistical mechanics and where $S_{E}[\phi]$ is the positive-definite Euclidean action. The mapping to imaginary time has been so far the unavoidable condition to set up any importance sampling numerical protocol to study the quantum fluctuations of fields. The goal of the present work is to present a conceptual framework and a method which goes beyond this limitation, allowing for a straightforward procedure to numerically sample quantum fluctuations of fields in Minkowskian spacetime. The use of imaginary time and Euclidean field theory forbids the representation on the lattice of any process or phenomenon intrinsically related to the causal structure of space-time, in particular all processes on the light cone. By definition the probability density $\exp(-S_{E}[\phi]/\hbar)$ works as an effective “equilibrium” measure for quantum fluctuations. Any importance-sampling protocol built from the Euclidean weight projects, for large lattice sizes, on the ground states of the corresponding Minkowskian theory. In fact, Monte Carlo simulations built from Euclidean field theory allow us to reproduce with extreme precision the physics of stable/equilibrium bound states of the strong interactions FH12 , whereas it has been so far much more problematic to reproduce metastable resonances with short lifetimes, like for instance tetraquark or pentaquark states MPR15 ; EPP17 , or the dynamics of scattering processes with a strong relativistic character, namely processes with a different number of degrees of freedom in the asymptotic initial and final states. It is for this reason that we believe it is of crucial interest the possibility to test numerically any new proposal for a quantum field theory formulation which allows first to define and then to study the dynamics of quantum fields fluctuations directly in Minkowski space-time.
@@
An interesting idea in this direction, namely the proposal of a functional approach to field theory which is well defined from the probabilistic point of view already in Lorentzian space-time, has been recently put forward by one of us and goes under the name of “symplectic quantization” GL21 ; G21 ; GLS24 . The first ingredient of this approach is to assume, for a given quantum field $\phi(x)$, with $x=(ct,{\bf x})$ a point in four-dimensional space-time, a dependence on an additional time parameter $\tau$:
$\phi(x)~\rightarrow~\phi(x,\tau),$ (1)
@@
which controls the continuous sequence of quantum fluctuations in each point of space-time. Theories with such an additional time parameter are not a novelty, the whole Parisi-Wu stochastic quantization approach being based on this idea PW81 ; DH87 .
Within the stochastic quantization approach, Euclidean multipoint correlation functions are obtained as a time average over the fictitous time parameter,
$\lim_{\tau\rightarrow\infty}\frac{1}{\tau}\int_{0}^{\tau}\phi(x_{1},s)\ldots\phi(x_{n},s)=\langle\phi(x_{1})\ldots\phi(x_{n})\rangle_{E},$ (2)
where on the right hand term of the above equation $\langle~\rangle_{E}$ denotes expectation with respect to the Euclidean weight and on the left hand term each $\phi(x_{i},s)$ represents the solution at point $x_{i}$ of an auxiliary Langevin dynamics of the kind
@@
where $\eta(x,\tau)$ represents is a zero mean white noise. It is a standard result in the theory of stochastic processes R96 that the stochastic dynamics of Eq. (3) allows to sample asymptotically the equilbrium distribution $\exp\left(-S_{E}[\phi]/\hbar\right)$. Shortly after the seminal idea of stochastic quantization, people realized that, on the basis of statistical ensemble equivalence, the Euclidean probability density of quantum fields can be sampled also by following the solutions of the Hamilton equations generated by a generalized Hamiltonian functional $\mathbb{H}_{E}[\pi(x),\phi(x)]$ of the kind
$\mathbb{H}_{E}[\pi(x),\phi(x)]=\mathbb{K}_{E}[\pi(x)]+S_{E}[\phi(x)],$ (4)
where $\pi(x,\tau)$ is a generalized momentum conjugated to the field $\phi(x,\tau)$ with respect to the flowing of the fictious time $\tau$. The idea of CR82 ; DFF83 , which goes under the name of “Microcanonical approach to quantum field theory”, is to achieve a sampling of the Euclidean probability of the fields by studying the following deterministic equations:
$\dot{\phi}(x,\tau)$ $=\frac{\delta}{\delta\pi(x)}\mathbb{H}_{E}[\pi,\phi]$
$\dot{\pi}(x,\tau)$ $=-\frac{\delta}{\delta\phi(x)}\mathbb{H}_{E}[\pi,\phi].$ (5)
The replacement of the stochastic dynamics in Eq. (3) with the deterministic one in Eq. (5) was solely motivated by its major computational efficiency in certain specific situations, for instance in the case of non-local bosonic actions obtained from the integration of fermionic variables MM94 , where the deterministic equations are more suited to parallel updates of the variables. As such, the role of the microcanonical ensemble built on the the conservation of the generalized Hamiltonian $\mathcal{H}_{E}[\pi(x),\phi(x)]$ was merely that of an alternative technique to sample $\exp\left(-S_{E}[\phi]/\hbar\right)$, where no physical interpretation was given neither to the additional time $\tau$ nor to the conjugated momenta $\pi(x,\tau)$. On the contrary, the key idea of symplectic quantization is to claim that the microcanonical approach to quantum field theory is something more fundamental and general, valid independently from its formal equivalence to the Euclidean field theory: this, as we are going to show, allows to sample quantum fluctuations directly in Minkowskian space-time. The logic of the following exposition will be therefore precisely the opposite of the one used to introduce the microcanonical approach to Euclidean quantum field theory, justified solely on its formal equivalence with the latter. First, without knowking which is the corresponding canonical ensemble, we claim the existence of a microcanonical ensemble built on a generalized Hamiltonian of the kind
$\mathbb{H}[\pi(x),\phi(x)]=\mathbb{K}[\pi(x)]-S[\phi(x)],$ (6)
where again we have a generalized kinetic energy term $\mathbb{K}[\pi(x)]$ and where this time the generalized potential is $S[\phi(x)]$, with $S[\phi(x)]$ the original Minkowskian action. We will show that the two point correlation function obtained by generating the quantum fluctuations of a weakly non-linear $\lambda\phi^{4}$ theory from the dynamics generated by $\mathbb{H}[\pi(x),\phi(x)]$ looks precisely like the standard Feynman propagator. It will be only after having checked that the deterministic dynamics generated by $\mathbb{H}[\pi(x),\phi(x)]$ allows to reproduce results in qualitative agreement to ordinary quantum field theory that we will find out which is the canonical weigth corresponding to our deterministic dynamics, obtaining it as the main result of the Sec. 7. In particular, in Sec. 7 we will show that in the thermodynamic limit the sampling of quantum fluctuation with our symplectic dynamics is equivalent to the sampling according to a canonical weight of the kind
@@
where the results of our simulations correspond to fixing $z=1$ in Eq. (7). Clearly, a probability density like the one in Eq. (7) is not equivalent to the complex amplitude which enters the Feynman path integral, but can be related to it by analytic continuation in the complex $z$ plane. Investigations in this direction are in progress and will be presented in a forthcoming paper explaining how the symplectic quantization approach works for a quantum particle in a harmonic potential, with particular emphasis on how the analytic continuation to complex $z$ plane of the above density, Eq. (7), must be handled GSG25 .
In synthesis, main idea of the symplectic quantization approach is the possibility to sample quantum fluctuation directly in Lorentzian space-time by means of a generalized microcanonical ensemble, built by adding the conjugated momenta with respect to the intrinsic time $\tau$ directly to the relativistic action with Minkowski signature, with no need of considering any sort of analytic continuation from real to immaginary time and therefore preserving the causal structure of space-time. This intuition has been strongly inspired from the evidence that in statistical mechanics there are physically important situations where only the microcanonical ensemble is well defined GILM21 ; GILM21b , namely physical phenomena which cannot be described within the canonical ensemble.
2 Symplectic Quantization: from dynamics to ensemble averages
Let us summarize here the main steps for the derivation of the symplectic quantization dynamics. First of all, inspired by the stochastic quantization approach PW81 ; DH87 , we assume that quantum fields $\phi(x,\tau)$ depend on an additional time variable $\tau$ which parametrizes the dynamics of quantum fluctuations in a given point of Minkowski space-time. Since for a relativistic quantum field theory the ambient space includes observer’s time, necessarily the intrinsic time $\tau$ must be a different variable, as thoroughly discussed in GL21 ; G21 . The symplectic quantization approach to field theory assumes, consistently with the existence of an intrinsic time $\tau$, the existence of conjugated momenta of the kind
@@ 2 Symplectic Quantization: from dynamics to ensemble averages
$\mathbb{L}[\phi,\dot{\phi}]=\int d^{d}x~\left[\frac{1}{2c_{s}^{2}}\dot{\phi}^{2}(x)+S[\phi]\right],$ (9)
@@
$\mathbb{H}[\phi,\pi]$ $=\frac{1}{2}\int d^{d}x~c_{s}^{2}~\pi^{2}(x)-S[\phi]$
@@
$=\int d^{d}x~\left[\frac{c_{s}^{2}}{2}\pi^{2}(x)+\frac{1}{2}\phi~\partial_{0}^{2}\phi-\sum_{i=1}^{d}\phi~\partial_{i}^{2}\phi+V[\phi]\right]$ (12)
@@
$\mathbb{H}[\phi,\pi]$ $=\mathbb{K}[\pi]+\mathbb{V}[\phi]$ (15)
with:
@@
$\mathbb{K}[\pi]$ $=\frac{1}{2}\int d^{d}x~\pi^{2}(x).$ (16)
@@
Having defined the above deterministic dynamics for the quantum fluctuations of the field $\phi(x,\tau)$, Eq. (14), one can then legitimately wonder how this functional formalism connects to the standard one, for instance to the standard Feynman path-integral formulation of quantum field theory. In first instance, in order to find out the connection between the dynamic approach of symplectic quantization and any functional formulation of field theory where the additional time $\tau$ is absent, one needs to find out which probability density $\rho[\phi(x)]$ corresponds to the dynamics. There are then two possibilities: $\rho[\phi(x)]$ might have either an integral or a punctual correspondence with quantum field theory. The integral correspondence is when the path-integral is recovered by means of an integral transformation, as it will be discussed below here, and the punctual correspondence is when, in a certain limit, one can directly show that $\rho[\phi(x)]\approx e^{iS[\phi]/\hbar}$. This last procedure is more subtle, because it requires an analytic continuation of the action and of its degrees of freedom in the complex plane, and will be discussed thoroughly in GSG25 ; GSG25_2 . But let us now go back to the relation between symplectic dynamics and probability densities, for which we make an ergodic hypothesis for the Hamiltonian dynamics in Eq. (13): if we assume that this dynamics samples at long time $\tau$ the constant generalized energy hypersurface with uniform probability GL21 ; G21 , then we can associate to the dynamics of Eq. (13) the following measure:
$\rho_{\text{micro}}[\phi(x)]=\frac{1}{\Omega[\mathcal{A}]}~\delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]\right),$ (18)
@@
$\Omega[\mathcal{A}]=\int\mathcal{D}\phi\mathcal{D}\pi~\delta\left(\mathcal{A}-\mathbb{H}[\phi,\pi]\right),$ (19)
@@
$\lim_{\Delta\tau\rightarrow\infty}\frac{1}{\Delta\tau}\int_{\tau_{0}}^{\Delta\tau}d\tau~\mathcal{O}[\phi(x,\tau)]=\int\mathcal{D}\phi\mathcal{D}\pi~\rho_{\text{micro}}[\phi(x)]~\mathcal{O}[\phi(x)],$ (21)
where $\tau_{0}$ is a large enough time for the system to have reached stationarity and “lost memory” of initial conditions. How to relate then the microcanonical partition function in Eq. (19) to the path integral? It is quite intuitive to understand that the two expression must be related by some sort of statistical ensemble change. The crucial point of this change of ensemble, as stressed already in GL21 , is that the microcanonical partition function $\Omega[\mathcal{A}]$ is built on the conservation of a non-positive quantity, the generalized Hamiltonian $\mathbb{H}[\phi,\pi]$. The latter, from the point of view of physical dimensions, is a relativistic action: it therefore takes both arbitrarily large positive and arbitrarily large negative values due to the negative sign in front of the coordinate-time derivative term in the second line of Eq. (12). The absence of positive definiteness for the generalized Hamiltonian $\mathbb{H}[\phi,\pi]$, which is the true relativistic signature of the theory, is what forbids a standard change of ensemble with a Laplace transform, that is customary in statistical mechanics when passing from microcanonical to canonical. The only integral transform which allows us to map formally the ensemble where $\mathbb{H}[\phi,\pi]$ is constrained to the one where it is free to fluctuate is the Fourier transform. It is by Fourier transforming the microcanonical partition function $\Omega[\mathcal{A}]$ that one obtains straightforwardly the Feynman path integral:
$\mathcal{Z}[u]$ $=\int_{-\infty}^{\infty}dA~e^{-iu\mathcal{A}}~\Omega[\mathcal{A}]=\int\mathcal{D}\phi~\mathcal{D}\pi~e^{-\frac{i}{2}u\int d^{d}x~\pi^{2}(x)+iuS[\phi]}=\mathcal{N}(u)\int\mathcal{D}\phi~e^{iuS[\phi]},$ (22)
@@
$\mathcal{Z}[\hbar]=\int_{-\infty}^{\infty}d\mathcal{A}~e^{-iA/\hbar}~\Omega[\mathcal{A}]\propto\int\mathcal{D}\phi~e^{\frac{i}{\hbar}S[\phi]}.$ (23)
The one above is to our knowledge the first derivation from first principles of the Feynman path-integral formula in the context of a more extended framework. We could say that this larger framework is the statistical mechanics of action-preserving systems, opposed to the statistical mechanics of energy-preserving systems, which is the standard one.
At this stage one can therefore legitimately wonder which is the relation between the probability density of the extended framework just introduced, i.e., the $\rho_{\text{micro}}[\phi(x)]$ of Eq. (18), and standard quantum field theory probability amplitudes. The key to this, at the present stage of development of this new approach, is precisely the Fourier transform of Eq. (23): what we expect from the microcanonical symplectic quantization ensemble is the possibility to sample disconnected correlation functions at fixed generalized action $\mathcal{A}$ and then, by Fourier transforming with respect to $\mathcal{A}$, obtain the original disconnected correlation functions of quantum field theory. We will see in Sec. 6 that remarkably good results for the two point correlation function can be obtained alredy in the fixed generalized action ensemble! We speak about disconnected correlation functions because, at the present stage of understanding, we are able to write down an explicit relation only between the generating functionals of disconnected correlations, respectively $\Omega[\mathcal{A}]$ for symplectic quantization and $\mathcal{Z}[\hbar]$ for quantum field theory. Clearly the new approach needs to be improved, for at least three main reasons. First, to recover disconnected correlation functions of quantum field theory by doing many deterministic simulations at different fixed values of action $\mathcal{A}$ and then try to Fourier transform the result looks a quite impractical protocol. We are presently working on an attempt to improve the correspondence between symplectic quantization and quantum field theory, but is still in progress GSG25 ; GSG25_2 . Second, the physical information of a quantum field theory is contained in the connected correlators, not the disconnected ones, so to rebuild the whole theory from this path seems a very long way. Last, but not least, as we find in the result of simulations in Sec. 5 and we will derive analytically in Sec. 7, the microcanonical density of Eq. (18) is ill-defined for a free theory, which looks quite problematic when compared to the importance and the success of perturbative expansions in quantum field theory.
Let us now come back for a moment on the reason why, differently from what is customary in the theory of statistical ensembles, symplectic quantization is related to quantum field theory by Fourier rathern than Laplace transforming. This comes from the fact that the microcanonical statistical ensemble of sympletic quantization is built on the conservation of a non positive-defined quantity, landmark of the relativistic nature of the theory, that implies the necessity of Fourier transforming and determines the fact that locally we can only access complex probability amplitudes and not real probabilities. From the perspective of symplectic quantization the replacement at the local level of probabilities with probability amplitudes is therefore a direct consequence of special relativity and a wise use of statistical ensembles. To better understand this statement let us consider an unrealistic situation where the symplectic action (generalized Hamiltonian) $\mathbb{H}[\phi,\pi]$ was positive definite. In this case one could change ensemble with Laplace rather than Fourier transform,
$\mathcal{Z}[\mu]=\int_{0}^{\infty}d\mathcal{A}~e^{-\mu\mathcal{A}}~\Omega[\mathcal{A}]\propto\int\mathcal{D}\phi~e^{\mu S[\phi]},$ (24)
@@
We have just shown how the standard path-integral formulation can be recovered, on the basis of an ergodicity assumption, from the symplectic quantization dynamics approach and which is the role played by $\hbar$ within this, let us say, change of ensemble. At the same time it is not only legitimate but also necessary to wonder if and how there is a quantization constraint involving $\hbar$ which can be imposed directly on the microcanonical ensemble of symplectic quantization. The indication coming from the stochastic quantization framework is that $\hbar$ must play a role analogous to that of temperature. Therefore, as suggested in GLS24 , we believe that the most natural assumption for the role of $\hbar$ in the symplectic quantization formalism is to be analogous to the microcanonical temperature:
@@
$\langle\pi(x)\pi(y)\rangle=\frac{\hbar}{2}~\delta^{(d)}(x-y),$ (26)
@@
$\langle\pi(x)\pi(y)\rangle=\lim_{\Delta\tau\rightarrow\infty}\frac{1}{\Delta\tau}\int_{\tau_{0}}^{\Delta\tau}d\tau~\pi(x,\tau)\pi(y,\tau)$ (27)
@@
$\langle\pi(x_{i})\pi(x_{j})\rangle=\frac{\hbar}{2}~\frac{\delta_{ij}}{a^{d}},$ (28)